# Understanding the Drag Torque in Common Envelope Evolution

### v1

@@ 4 RESULTS
Figure 2: The $z$-component of torque on the binary system about the particle CM (left axis). Top: Shown is the torque (i) measured directly from the simulation (black), (ii) including only contributions out to the contour $\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$ (magenta), (iii) reconstituted using equation (10) with $\overline{\rho}$, $\Delta\phi$ and $\widetilde{B}/\widetilde{A}$ measured from the simulation (cyan), (iv) reconstituted using equation (24) for a co-rotating spheroid with $\langle\Delta\phi\rangle=14.9^{\circ}$ and $\langle\widetilde{B}/\widetilde{A}\rangle=0.654$ (orange). The orbital separation of the particles, $a$, is shown on the right axis (dashed red). The inset shows a zoom-in of the torque at late times. Bottom: The $z$-component of the torque (i) measured directly from the simulation (same as in the top panel, black), (ii) calculated from equation (22) with $c_{\mathrm{s,0}}$ taken as the mean sound speed inside the surface $\rho=\rho_{\mathrm{c}}$ and $\rho_{0}$ taken as the density on this surface (dashed magenta), and (iii) the same but now $\rho_{0}$ is taken to be $0.44$ times the value on the surface (solid magenta). The Mach number of each particle is shown on the right axis (green, dotted for $\mathcal{M}_{1}$ and dashed for $\mathcal{M}_{2}$; see equation 20).
[…]
@@
Figure 7: From left to right, columns show snapshots in the orbital plane at times $t=138.9$, $208.3$, and $277.8$ days. Primary core (left) and companion (right) softening spheres are shown in black. Rows from top to bottom are: (i) Mass density ($\rho$, thick white contour for $\rho=\rho_{\mathrm{c}}$, and thin grey contours showing $\rho=4\rho_{\mathrm{c}},2\rho_{\mathrm{c}},0.5\rho_{\mathrm{c}}\,\text{and}\,0.25\rho_{\mathrm{c}}$); (ii) The Mach number of the perturbers $(\mathcal{M}_{1}+\mathcal{M}_{2})/2$ (blue being the subsonic region and red being the supersonic region); (iii) The difference over the sum of the angular speeds of the perturbers and gas $(\Omega-\Omega_{\mathrm{gas}})/(\Omega+\Omega_{\mathrm{gas}})$ (red being the region dominated by $\Omega$, blue being the region dominated by $\Omega_{\mathrm{gas}}$); (iv) The torque density on both particles. Black and white contours respectively show the positively and negatively contributing regions to the $\phi$-component of the drag. The highest contour levels near the particles are ±$\pm 10^{12}\,{\rm dyn\,cm^{-2}}$, and contours are also plotted for ±$\pm 10^{11}\,{\rm dyn\,cm^{-2}}$, ±$\pm 10^{10}\,{\rm dyn\,cm^{-2}}$, and so on. Movies starting from $t=115.7$ days to the end of the simulation are available at https://doi.org/10.5281/zenodo.15718017.
@@ 4.1 Orbit
The evolution of the distance $a$ between the AGB core particle and companion is shown in Figure 1, and the orbit of the particles is shown in the inset. These results are similar to those in other CE simulations (e.g. Ohlmann et al., 2016a; Chamandy et al., 2019b, 2020).
@@ 4.2 Torque exerted on the particles by the gas
In the top panel of Figure 2, we show the evolution of the $z$-component of the torque applied on the particles about their CM (left axis, solid black), from $t\approx 115\,{\rm d}$ onward. The evolution of the orbital separation is shown for comparison (right axis, dashed). The drastic reduction of the torque with time is consistent with previous studies involving different initial binary parameter values (Chamandy et al., 2019a, 2020).
In order to model this torque, we first need to determine which part of the gas contributes significantly to it. This is done by integrating the torque out to various trial values of $\rho/\rho_{\mathrm{max}}$, with $\rho_{\mathrm{max}}$ the maximum density in a slice through the orbital plane, for a given snapshot (throughout the simulation, the density maximum occurs near the position of the AGB core particle). It is found that using the trial value $\rho/\rho_{\mathrm{max}}=0.006$ leads to a torque that deviates the least from the actual torque. In the top panel of Figure 2, it can be seen that the torque exerted by gas inside the contour $\rho/\rho_{\mathrm{max}}=0.006$ (magenta) matches closely with that exerted by all the gas (black). Henceforth, we refer to this as the threshold or critical density $\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$.
@@ 4.3 Fitting an ellipsoid to the threshold density contour
To apply the uniform density ellipsoid model of Section 3.1, it seems natural to use the ellipsoid that best approximates the contour at the threshold density $\rho=0.006\rho_{\mathrm{max}}$. To simplify the fitting procedure, we choose to fit two ellipses in perpendicular planes (which uniquely define an ellipsoid) rather than directly fitting an ellipsoid. We first fit the $\rho=\rho_{\mathrm{c}}$ contour in the orbital plane with a co-planar ellipse, treating the aspect ratio $B/A$ and phase shift $\Delta\phi$ as fit parameters. We then fit the $\rho=\rho_{\mathrm{c}}$ contour in the plane perpendicular to the orbital plane that contains the ellipse major axis (i.e. the plane rotated by $\Delta\phi$ from the line joining the particles), setting the length of the semi-major axis of this ellipse also equal to $A$ but fitting for the semi-minor axis $C$. These two ellipses define an ellipsoid with major axis and one other axis in the orbital plane, and the third axis perpendicular to the orbital plane. The top two panels of Figure 3 show an example of such a fit, at the snapshot $t=188.7\,{\rm d}$, where the phase shift ($\Delta\phi$) is equal to $14.7^{\circ}$ (which also happens to be almost equal to the mean value of $\Delta\phi$ over the time domain analyzed, as discussed below). The blue ellipses show the fits to the $\rho=\rho_{\mathrm{c}}$ blue contour in the orbital plane (left) and orthogonal plane containing the ellipse major axis (right).
Our interest is in modeling the slow spiral-in phase, and the models cannot be applied right from $t=0$. To determine the appropriate time domain to apply the ellipsoid model, we qualitatively judged when the density contours in the orbital plane resemble ellipses, which led to the estimate $t>125\,{\rm d}$. This choice is supported by quantitative analysis shown in Figure 4, where we plot the root mean square error (RMSE) of our ellipsoidal fit, normalized to be between 00 and $1$. It is seen that the RMSE decreases below $0.1$ after $t=125\,{\rm d}$ and remains below this value thereafter.
@@ 4.4 Mass evolution
In Figure 6 (left axis), we show the ratio of the mass of the gas enclosed by the $\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$ surface to the mass $m=M_{\mathrm{1,c}}+M_{2}$ of the particles. This ratio decreases with time as the particles inspiral and is always $\ll 1$. This is the regime for which E04 applied their uniform ellipsoid model. The mean density inside the $\rho=\rho_{\mathrm{c}}$ surface is shown in blue (right axis). If we compare $\overline{\rho}$ with $A/a$ (red curve in Figure 5), we can observe that the relationship between these quantities is complicated, despite the fact that $A/a$ enters into the model only insofar as it is used for determining $\overline{\rho}$.
@@ 4.5 To what extent can the uniform spheroid model reproduce the torque?
where $\widetilde{\alpha}=0.457$ and $\widetilde{\beta}=0.771$ are the values of $\alpha$ and $\beta$ when $\widetilde{B}/\widetilde{A}$ is replaced by $\langle\widetilde{B}/\widetilde{A}\rangle$. Equation (24) was also used by E04 to model equal-mass inspiralling BSMBHs (see Section 1). We find that this approximation, shown as an orange line in the top panel of Figure 2, aligns quite well with the actual torque (black), though it introduces a phase shift of about a half period. The phase difference between the model and the actual torque can be explained by the direct dependence of equation (24) on the square of the separation $a$, which causes the modeled torque to be in phase with the separation.
@@ 4.6 To what extent can solutions of the idealized binary perturbers problem reproduce the torque?
In this section, we compare the torque evolution in the simulation with that predicted by the K08 model, i.e., equation (19). The K08 model assumes that the unperturbed medium is uniform, and in order to apply this model, one needs to specify the unperturbed density $\rho_{0}$ and unperturbed sound speed $c_{\mathrm{s,0}}$. It is not clear how this should be done, so we tried different choices. One possible choice is to set $\rho_{0}(t)$ to the value of the density on the surface that bounds the region that contributes significantly to the torque, $\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$, and $c_{\mathrm{s,0}}(t)$ to the mean sound speed interior to this surface. In practice, we introduce an adjustable scale factor multiplying $\rho_{0}(t)$, which is expected to be of order unity, so that we finally take
@@
The K08 torque on the perturbers is then calculated using equation (22), and is shown by a magenta line in the bottom panel of Figure 2. The modeled torque is seen to agree quite well with the simulation from $t=270\,{\rm d}$ onward. For $t<250\,{\rm d}$, the K08 torque is too large by a factor of order unity; after multiplying the modeled torque by $0.44$, it reproduces the simulation torque remarkably well, including (to varying degrees) the magnitude, amplitude and temporal phase evolution.
@@
We also plot the Mach number of each particle using the right axis of the bottom panel of Figure 2. The Mach number of the AGB core particle, $\mathcal{M}_{1}$, is represented by a dotted green line and that of the companion particle, $\mathcal{M}_{2}$, by a dashed green line. They can be seen to vary on the orbital period and also on longer timescales, but their time averages are very similar. Their values are restricted to the range 0.50.50.50.50.70.70.70.7, so the particles’ azimuthal motion is subsonic.
@@
For completeness we now mention alternative choices for specifying $\rho_{0}(t)$ and $c_{\mathrm{s,0}}(t)$. As a second approach, we tried setting $c_{\mathrm{s,0}}(t)$ to its value on the $\rho=\rho_{\mathrm{c}}$ surface, which leads to a larger Mach number and thus values of the torque that are about a factor of two larger than above, so about five times larger than that measured from the simulation. A third approach we tried was to take both $\rho_{0}(t)$ and $c_{\mathrm{s,0}}(t)$ to be equal to the average values inside the $\rho=\rho_{\mathrm{c}}$ surface. This produces results similar to the first method ($\rho_{0}$ equal to the value at the surface), but the torque is similar in magnitude to what is obtained in the second approach. Fourth, we considered using the original profile of the AGB star by setting $\rho(t)=\rho(r)$ with $r=a(t)$, and likewise for the sound speed. This would be convenient if it worked, because then one would only need to know the initial conditions. However, this choice produces torque values that are higher than those measured in the simulation by about an order of magnitude, and a torque evolution that does not resemble the simulation even after dividing by a constant numerical factor. The failure of this fourth approach suggests that it is not straightforward to model accurately the torque at late times using the initial stellar profile and K08 model alone, at least not for companions whose masses are appreciable compared to that of the giant’s core.
@@
To summarize, the K08 model reproduces the simulation remarkably well between $t\approx 125\,{\rm d}$ and $t\approx 250\,{\rm d}$ when one adopts $\rho_{0}=0.44\rho_{\mathrm{c}}(t)$ and $c_{\mathrm{s,0}}=\overline{c}_{\mathrm{s}}(t)$, whereas at late times choosing $\rho_{0}=\rho_{\mathrm{c}}(t)$ produces closer agreement. This is discussed further in Section 5.2.1.

### v3

@@ 4 RESULTS
Figure 2: The $z$-component of torque on the binary system about the particle CM (left axis). Top: Shown is the torque (i) measured directly from the simulation (black), (ii) including only contributions out to the contour $\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$ (magenta), (iii) reconstituted using equation (10) with $\overline{\rho}$, $\Delta\phi$ and $\widetilde{B}/\widetilde{A}$ measured from the simulation (cyan), (iv) reconstituted using equation (24) for a co-rotating spheroid with $\langle\Delta\phi\rangle=14.9^{\circ}$ and $\langle\widetilde{B}/\widetilde{A}\rangle=0.654$ (orange). The orbital separation of the particles, $a$, is shown on the right axis (dashed red). The inset shows a zoom-in of the torque at late times. Bottom: The $z$-component of the torque (i) measured directly from the simulation (same as in the top panel, black), (ii) calculated from equation (22) with $c_{\mathrm{s,0}}$ taken as the mean sound speed $\overline{c}_{\mathrm{s}}$ inside the surface $\rho=\rho_{\mathrm{c}}$ and $\rho_{0}$ taken as the density on this surface (dashed magenta), and (iii) the same but now $\rho_{0}$ is taken to be $0.44$ times the value on the surface (solid magenta). The particle Mach number $\mathcal{M}_{\mathrm{p}}$, obtained by dividing the particle tangential speed in the particle CM frame by $\overline{c}_{\mathrm{s}}$, is shown on the right axis in solid green.
Figure 3: Top left: Density contours at $\rho=0.01\rho_{\mathrm{max}}(t)$, $0.006\rho_{\mathrm{max}}(t)$, and $0.005\rho_{\mathrm{max}}(t)$ in the orbital plane at the time $t=188.7\,{\rm d}$, with ellipse fitted to the contour $\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$, which was found to enclose effectively all of the gas producing significant torque (see also Figure 2). The ellipse is phase-shifted by an angle $\Delta\phi$ with respect to the axis that passes through the particles. Top right: Similar to the top-left panel but now showing the plane perpendicular to the orbital plane and rotated clockwise by the angle $\Delta\phi$ about the orbital axis, shown by the dashed line in the top left panel. The length of the ellipse major axis is set equal to that in the orbital plane, but the length of the minor axis is allowed to differ. Bottom left: Adapted from figure 1 of Kim+08, showing the steady state for $\mathcal{M}_{\mathrm{p}}=0.6$ sliced through the orbital plane in their idealized double-perturber model. The black circle shows the orbit of the point masses that perturb the background density. Colour shows the density contrast $\log\mathcal{D}$, where $\mathcal{D}=(c_{\mathrm{s,0}}^{2}a/Gm)\lambda$ with $m$ the binary mass and $\lambda=(\rho-\rho_{0})/\rho_{0}$. Overplotted for comparison is the time-averaged best fit ellipse in the orbital plane from our simulation, with parameter values noted on the plot (see also Figure 5). Bottom right: Similar to the bottom left panel but now showing $\log\mathcal{D}$ for our simulation, at the same time as the top row, when $\rho_{0}=0.44\rho_{\mathrm{c}}=3.16\times 10^{-6}\,{\rm g\,cm^{-3}}$, and $c_{\mathrm{s,0}}=\overline{c}_{\mathrm{s}}=93.0\,{\rm km\,s^{-1}}$. The region outside $0.44\rho_{\mathrm{c}}$ has negative values of $\mathcal{D}$ and is coloured grey. The contour $\rho=\rho_{\mathrm{c}}$ is plotted in yellow. The white contours near the softening radius (black circles, AGB core on the left and companion on the right) show contours of $\log\mathcal{D}=1.2\,\mathrm{and}\,0.8$, while the contours outside $\rho=\rho_{\mathrm{c}}$ show $\log\mathcal{D}=-0.4,-0.8\,\mathrm{and}\,-1.2$.
@@
Figure 7: From left to right, columns show snapshots in the orbital plane at times $t=138.9$, $208.3$, and $277.8$ days. Primary core (left) and companion (right) softening spheres are shown in black or yellow in the bottom row. Rows from top to bottom are: (i) Mass density ($\rho$, thick white contour for $\rho=\rho_{\mathrm{c}}$, and thin grey contours showing $\rho=4\rho_{\mathrm{c}},2\rho_{\mathrm{c}},0.5\rho_{\mathrm{c}}\,\text{and}\,0.25\rho_{\mathrm{c}}$); (ii) The local Mach number of the particles $V_{\phi}/c_{\mathrm{s}}$, where $V_{\phi}$ is the $\phi$-component of the particle velocity in the particle centre of mass frame ($V_{1,\phi}=V_{2,\phi}$) and $c_{\mathrm{s}}$ is the sound speed, which depends on position (blue denotes the subsonic region and red the supersonic region); (iii) The difference over the sum of the angular speeds of the perturbers and gas $(\Omega-\Omega_{\mathrm{gas}})/(\Omega+\Omega_{\mathrm{gas}})$ (red being the region dominated by $\Omega$, blue being the region dominated by $\Omega_{\mathrm{gas}}$); (iv) The torque density on both particles. Black and white contours respectively show the negatively and positively contributing regions to the $\phi$-component of the drag. The highest contour levels near the particles are $\pm 10^{12}\,{\rm dyn\,cm^{-2}}$, and contours are also plotted for $\pm 10^{11}\,{\rm dyn\,cm^{-2}}$, $\pm 10^{10}\,{\rm dyn\,cm^{-2}}$, and so on. Movies starting from $t=115.7$ days to the end of the simulation are available at https://doi.org/10.5281/zenodo.17575148.
@@ 4.1 Orbit
The evolution of the distance $a$ between the AGB core particle and companion is shown in Figure 1, and the orbit of the particles is shown in the inset. These results are similar to those in other CE simulations (e.g. Ohlmann+16a; Chamandy+19a; Chamandy+20).
@@ 4.2 Torque exerted on the particles by the gas
In the top panel of Figure 2, we show the evolution of the $z$-component of the torque applied on the particles about their CM (left axis, solid black), from $t\approx 115\,{\rm d}$ onward. The evolution of the orbital separation is shown for comparison (right axis, dashed). The drastic reduction of the torque with time is consistent with previous studies involving different initial binary parameter values (Chamandy+19b; Chamandy+20).
In order to model this torque, we first need to determine which part of the gas contributes significantly to it. This is done by integrating the torque out to various trial values of $\rho/\rho_{\mathrm{max}}$, with $\rho_{\mathrm{max}}$ the maximum density in a slice through the orbital plane, for a given snapshot (throughout the simulation, the density maximum occurs near the position of the AGB core particle). It is found that using the trial value $\rho/\rho_{\mathrm{max}}=0.006$ leads to a torque that deviates the least from the actual torque. In the top panel of Figure 2, it can be seen that the torque exerted by gas inside the contour $\rho/\rho_{\mathrm{max}}=0.006$ (magenta) matches closely with that exerted by all the gas (black). Henceforth, we refer to this as the threshold or critical density $\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$. Further justification for the value $0.006$ is provided in B.
However, it is important to note that $\rho_{\mathrm{max}}(t)$ depends on the choices involved in modifying the initial mesa profile and softening the gravity around the AGB core particle (Section 2). Therefore, the value of $\rho_{\mathrm{max}}$ (and by extension, the ratio $\rho_{\mathrm{c}}/\rho_{\mathrm{max}}$) has limited physical significance. This is not a major concern because the role of $\rho_{\mathrm{max}}$ is simply to provide a convenient normalization for the density. We shall see below that the linear spatial scale of the contour $\rho=\rho_{\mathrm{c}}(t)$ is roughly proportional to the core-companion separation $a(t)$ and that the morphological evolution is approximately self-similar, as also seen in Escala+04. This self-similarity helps to explain why setting $\rho_{\mathrm{c}}$ to be a fixed fraction of $\rho_{\mathrm{max}}$ turns out to be a fruitful approach. A choice of reference density other than $\rho_{\mathrm{max}}$, e.g., that at the particle CM or averaged over the circle with radius $a/2$ centred on the particle CM, could perhaps be used instead, and we plan to explore this possibility in future work.
@@ 4.3 Fitting an ellipsoid to the threshold density contour
To apply the uniform density ellipsoid model of Section 3.1, it seems natural to use the ellipsoid that best approximates the contour at the threshold density $\rho=0.006\rho_{\mathrm{max}}$. To simplify the fitting procedure, we choose to fit two ellipses in perpendicular planes (which uniquely define an ellipsoid) rather than directly fitting an ellipsoid. We first fit the $\rho=\rho_{\mathrm{c}}$ contour in the orbital plane with a co-planar ellipse, treating the semi-major axis $A$, aspect ratio $B/A$ and phase shift $\Delta\phi$ as fit parameters. We then fit the $\rho=\rho_{\mathrm{c}}$ contour in the plane perpendicular to the orbital plane that contains the ellipse major axis (i.e. the plane rotated by $\Delta\phi$ from the line joining the particles), setting the length of the semi-major axis of this ellipse also equal to $A$ but fitting for the semi-minor axis $C$. These two ellipses define an ellipsoid with major axis and one other axis in the orbital plane, and the third axis perpendicular to the orbital plane. The top two panels of Figure 3 show an example of such a fit, at the snapshot $t=188.7\,{\rm d}$, where the phase shift ($\Delta\phi$) is equal to $14.7^{\circ}$ (which also happens to be almost equal to the mean value of $\Delta\phi$ over the time domain analyzed, as discussed below). The blue ellipses show the fits to the $\rho=\rho_{\mathrm{c}}$ blue contour in the orbital plane (left) and orthogonal plane containing the ellipse major axis (right).
Our interest is in modeling the slow spiral-in phase, so the models cannot be applied right from $t=0$. To determine the appropriate time domain to apply the ellipsoid model, we qualitatively judged when the density contours in the orbital plane resemble ellipses, which led to the estimate $t>125\,{\rm d}$. This choice is supported by quantitative analysis shown in Figure 4, where we plot the root mean square error (RMSE) of our ellipsoidal fit, normalized to be between 0 and $1$. It is seen that the RMSE decreases below $0.1$ after $t=125\,{\rm d}$ and remains below this value thereafter.
@@ 4.4 Mass evolution
In Figure 6 (left axis), we show the ratio of the mass of the gas enclosed by the $\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$ surface to the mass $m=M_{\mathrm{1,c}}+M_{2}$ of the particles. This ratio decreases with time as the particles inspiral and is always $\ll 1$. This is the regime for which Escala+04 applied their uniform ellipsoid model. The mean density inside the $\rho=\rho_{\mathrm{c}}$ surface is shown in blue (right axis). The mean density $\overline{\rho}$ is roughly constant between $t=125\,{\rm d}$ and $t=270\,{\rm d}$, but decreases sharply thereafter. This sudden decrease may be an artefact caused by the ratio $r_{\mathrm{soft}}/a$ becoming too large, which implies a larger fractional volume inside which the gravitational potential is underestimated (see also Section 5.2.1).
@@ 4.5 To what extent can the uniform spheroid model reproduce the torque?
where $\widetilde{\alpha}=0.457$ and $\widetilde{\beta}=0.771$ are the values of $\alpha$ and $\beta$ when $\widetilde{B}/\widetilde{A}$ is replaced by $\langle\widetilde{B}/\widetilde{A}\rangle$. Equation (24) was also used by Escala+04 to model equal-mass inspiralling BSMBHs (see Section 1). We find that this approximation, shown as an orange line in the top panel of Figure 2, aligns quite well with the actual torque (black), though it introduces a phase shift of about a half period. The phase difference between the model and the actual torque can be explained by the direct dependence of equation (24) on the square of the separation $a$, which causes the modeled torque to be in phase with the separation.
@@ 4.6 To what extent can solutions of the idealized binary perturbers problem reproduce the torque?
In this section, we compare the torque evolution in the simulation with that predicted by the Kim+08 model, i.e., equation (19). The Kim+08 model assumes that the unperturbed medium is uniform, and in order to apply this model, one needs to specify the unperturbed density $\rho_{0}$ and unperturbed sound speed $c_{\mathrm{s,0}}$. It is not clear how this should be done, so we tried different choices. One possible choice is to set $\rho_{0}(t)$ to the value of the density on the surface that bounds the region that contributes significantly to the torque, $\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}(t)$, and $c_{\mathrm{s,0}}(t)$ to the mean sound speed interior to this surface. In practice, we introduce an adjustable scale factor multiplying $\rho_{0}(t)$, which is expected to be of order unity, so that we finally take
@@
The Kim+08 torque on the perturbers is then calculated using equation (22), and is shown by magenta lines in the bottom panel of Figure 2, dashed for $\xi=1$ and solid for $\xi=0.44$. The modeled torque using $\xi=1$ is seen to agree quite well with the simulation from $t=270\,{\rm d}$ onward, but this may be a coincidence since numerical effects may cause the torque to be overestimated in the simulation at late times (see Section 5.2.1). For $t<250\,{\rm d}$, the Kim+08 torque obtained using $\xi=1$ is too large by a factor of order unity; after multiplying the modeled torque by $\xi=0.44$, the model reproduces the simulation torque remarkably well, including (to varying degrees) the magnitude, amplitude and temporal phase evolution.
@@
We also plot the particle Mach number $\mathcal{M}_{\mathrm{p}}$, equal to the particle tangential speed in the particle CM frame divided by the volume-averaged mean sound speed $\overline{c}_{\mathrm{s}}$ averaged inside the contour $\rho=\rho_{\mathrm{c}}$, in green using the right axis of the bottom panel of Figure 2. The value of $\mathcal{M}_{\mathrm{p}}$ can be seen to vary on the orbital period and also on longer timescales, but is restricted to the range 0.50.50.70.7, so the particles’ azimuthal motion is subsonic.
@@
For completeness we now mention alternative choices for specifying $\rho_{0}(t)$ and $c_{\mathrm{s,0}}(t)$. As a second approach, we tried setting $c_{\mathrm{s,0}}(t)$ to the value of $c_{\mathrm{s}}$ on the $\rho=\rho_{\mathrm{c}}$ surface, which leads to a larger Mach number and values of the torque that are about a factor of two larger than above. A third approach we tried was to take both $\rho_{0}(t)$ and $c_{\mathrm{s,0}}(t)$ to be equal to the average values inside the $\rho=\rho_{\mathrm{c}}$ surface. This produces results more similar to the chosen method (first approach) but the torque is similar in magnitude to what is obtained in the second approach. From the bottom left panel of Fig. 3, one can see that the density at the contour $\rho=\rho_{\mathrm{c}}$ exceeds the background density. Thus, to the extent that our simulation lines up with the Kim+08 model, it makes sense that the best-fit value of $\rho_{0}$ for reproducing the torque is smaller than $\rho_{\mathrm{c}}$.
Fourth, we considered using the original profile of the AGB star by setting $\rho(t)=\rho(r)$ with $r=a(t)$, and likewise for the sound speed. This would be convenient if it worked, because then one would only need to know the initial conditions. However, this choice produces torque values that are higher than those measured in the simulation by about an order of magnitude, and a torque evolution that does not resemble the simulation even after dividing by a constant numerical factor. The failure of this fourth approach suggests that it is not straightforward to model accurately the torque at late times using the initial stellar profile and Kim+08 model alone, at least not for companions whose masses are appreciable compared to that of the giant’s core. This is perhaps not surprising because transfer of orbital energy and angular momentum to the envelope causes the profiles of density and sound speed to be drastically affected, as shown in Figs. 9 and 10 of C, and, as we argue in Section 5.2.2, the system loses memory of its previous state after about one sound-crossing time.
@@
To summarize, the Kim+08 model reproduces the simulation remarkably well between $t\approx 125\,{\rm d}$ and $t\approx 250\,{\rm d}$ when one adopts $\rho_{0}=0.44\rho_{\mathrm{c}}(t)$ and $c_{\mathrm{s,0}}=\overline{c}_{\mathrm{s}}(t)$, whereas at late times choosing $\rho_{0}=\rho_{\mathrm{c}}(t)$ produces closer agreement. This is discussed further in Section 5.2.1.